# CALM-PDE: Continuous and Adaptive Convolutions for Latent Space Modeling of Time-dependent PDEs

### v1

@@ 1 Introduction
Many scientific problems, such as climate modeling and fluid mechanics, rely on simulating physical systems, often involving solving spatio-temporal Partial Differential Equations (PDEs). In recent years, Machine Learning (ML) models have been successfully used to approximate the solution of PDEs (Lu et al., 2019; Li et al., 2020b; Brandstetter et al., 2022; Chen & Wu, 2024), offering several advantages over classical numerical PDE solvers. For instance, ML models offer a data-driven approach that is applicable even if the underlying physics is (partially) unknown, can generate solutions more efficiently (Li et al., 2020a; Tompson et al., 2022), and are inherently differentiable by design, which is often not the case for numerical solvers (Takamoto et al., 2022).
@@
Practical applications usually require a densely discretized spatial domain $\Omega$, leading to more than 1M spatial points per timestep. For example, ML-based weather forecasting models typically operate on a spatial domain of $720\times 1440$ points or pixels (Pathak et al., 2022). Learning the solution of a time-dependent PDE directly in the physical domain $\Omega$ rather than in a compressed latent space can result in high memory and computational costs. Consequently, several architectures for reduced-order PDE-solving have been introduced (Wu et al., 2022; Alkin et al., 2024; Serrano et al., 2024; Wang & Wang, 2024). These models adopt the encode-process-decode paradigm (Sanchez-Gonzalez et al., 2020), wherein the physical domain is encoded into a compact latent space, its dynamics are evolved via a processor, and the output is decoded back to the original domain. Many real-world applications involve geometries with irregularly sampled spatial domains, necessitating discretization-agnostic models to effectively process such data. Unfortunately, existing approaches mainly utilize Convolutional Neural Networks (CNNs), which necessitate a regular spatial discretization, or memory-intensive attention-based mechanisms (Vaswani et al., 2017) to enable the encoding and decoding of arbitrarily discretized spatial domains.
@@
Motivated by these considerations, we introduce CALM-PDE (Continuous and Adaptive Convolutions for Latent Space Modeling of Time-dependent PDEs), a novel architecture for reduced-order modeling of arbitrarily discretized PDEs. CALM-PDE compresses the spatial domain into a fixed latent representation with an encoder that builds on parametric continuous convolutional neural networks (Wang et al., 2018) and adaptively learns where to apply the convolution operator (learnable query points). This enables CALM-PDE to selectively sample more densely in important regions of the spatial domain, such as near complex solid boundaries, while allocating fewer points to smoother regions. Furthermore, we incorporate a locality inductive bias into the kernel function, consisting of an epsilon neighborhood and a distance-weighting, to enhance computational efficiency and to facilitate the learning of local patterns. An autoregressive model computes the temporal evolution completely in the latent space (latent time-stepping), which can be interpreted as Neural Ordinary Differential Equation (NODE; Chen et al. (2019)). The discretization-agnostic decoder, which builds on similar layers as the encoder, enables querying the output solution at arbitrary spatial locations. An overview of the CALM-PDE framework is shown in Figure 1.
@@
  • We propose a new model that solves arbitrarily discretized PDEs in a fixed latent space.
@@ 2 Problem Definition
In the following section, we formally introduce the problem of solving PDEs. We refer to Appendix B for information about the dataset and training objective to train neural surrogates for PDE-solving.
@@
where $\boldsymbol{u}:[0,T]\times\Omega\rightarrow\mathbb{R}^{N_{c}}$ represents the solution function of the PDE that satisfies the Initial Condition (IC) $\boldsymbol{u}(0,\boldsymbol{\omega})$ for $t=0$ and the boundary conditions $\mathcal{B}[\boldsymbol{u}](t,\boldsymbol{\omega})$ if $\boldsymbol{\omega}$ is on the boundary $\partial\Omega$ of the domain $\Omega$. $N_{c}$ denotes the number of output channels or field variables of the PDE. Solving a PDE involves computing (an approximation of) the function $u$ that satisfies Equation 1.
@@ 3 Background and Preliminaries
We briefly introduce deep parametric continuous convolutional neural networks (Wang et al., 2018). For comparison, discrete convolution and discrete CNNs are explained in Appendix C.
@@
[…]
@@
where $\mathtt{RF}(a_{j})$ outputs the indices for the $M$ points that lie in the receptive field of a query point $a_{j}$. The receptive field can be constructed by only considering the K-nearest neighbors of the point $a_{j}$ or by considering the points in an epsilon neighborhood. Setting $M:=N$ means that the receptive field is not limited. Appendix D shows a visualization of continuous convolution. To generalize the layer from 1D to d-D, only the input dimension of the kernel function has to be adapted.
@@ 4.1 Learning Continuous Convolutions with Adaptive Query Points
We follow the formulation of a deep parametric continuous convolutional layer in Equation 4 introduced by Wang et al. (2018) and propose the following improvements for encoding and decoding PDE solutions, which yield the CALM layer used for the encoder and decoder.
@@
We use a 2-layer MLP that takes $a-\alpha$ as input to parameterize the continuous kernel function $k_{i,o}(a-\alpha)$. The translation $a-\alpha$ is encoded with Random Fourier Features ($\mathtt{RFF}$; Li et al. (2021)) to allow the kernel function to be less smooth (i.e., a small change in the input translation can lead to large output changes in the weight) which could be beneficial for high-frequency content where slightly different translation vectors should result in a large difference. Thus, the parametrized kernel function is given as $k_{i,o}(a-\alpha)=\mathtt{MLP}\bigl{(}\mathtt{RFF}(a-\alpha)\bigl{)}_{i,o}$ where $i$ is the $i^{th}$ input channel and $o$ the $o^{th}$ output channel.
@@
[…]
for the kernel function. This emphasizes input points closer to the query point within the epsilon neighborhood. The $\mathtt{softmax}$ also serves as a normalization factor for the Monte Carlo integration. The combination of the limited receptive field and distance weighting is a locality inductive bias, which is helpful because local patterns are often more important in PDE-solving (Chen & Wu, 2024). Part (i) of Figure 2 shows the receptive field and the distance weighting within the receptive field.
@@
In continuous convolution, the input and output points do not have to be the same. This means that downsampling and upsampling are inherently supported, in contrast to CNNs, where strided convolution and variants are needed for downsampling. This allows us to reduce the number of points for encoding and to increase the number of points for decoding. We propose to make the query points for both encoding and decoding learnable. This allows the model to sample more densely in regions with important characteristics (e.g., a region that contains turbulence or an obstacle). The query positions are initially uniformly sampled from the domain $\Omega$. We assume that $\Omega$ is normalized in a range $[0,1)$ which yields $A=\{a_{j}\}_{j=1}^{K}=\{(x_{j},y_{j})\}_{j=1}^{K}$ with $x_{j},y_{j}\sim\mathcal{U}(0,1)$ for the initialization of the query points in 2D. Alternatively, the query points can be initialized by sampling $K$ query points from the underlying mesh. We call this initialization method of the query points as “mesh prior”. During the training, the query points are moved with a variant of stochastic gradient descent such as Adam (Kingma & Ba, 2017) as $(x_{j},y_{j})^{\top}\leftarrow(x_{j},y_{j})^{\top}-\eta\frac{\partial\mathcal{L}_{\boldsymbol{\theta}}}{\partial(x_{j},y_{j})^{\top}}\quad\forall j$ where $\eta$ denotes the learning rate and $\mathcal{L}_{\boldsymbol{\theta}}$ the cost function. The kernel function provides feedback in which direction to move, and the $\mathtt{softmax}$ function in Equation 5 avoids hard cuts caused by the learnable query points and epsilon neighborhood by weighting the input points on the boundary of the epsilon neighborhood lower. The learnable query points are illustrated in part (ii) of Figure 2.
@@
We allow each query point to have a customized filter kernel (i.e., different kernel weights for the same or similar translation vectors). This way, there can be multiple query points at the same location, each paying attention to different features. This could be helpful for simulations that contain a geometry or obstacles (e.g., many query points close to the obstacle, but each query learns different characteristics). The kernel modulation is done by scaling and shifting the intermediate representations (Perez et al., 2018) of the MLP in Equation 5. Each query point has a location as well as scaling and shifting parameters. We denote the modulated kernel function as
@@
where we exclude the distance weighting for the sake of simplicity. $\boldsymbol{W}$ denotes a weight matrix with a suitable shape, $\sigma$ is a non-linearity, and $\boldsymbol{\gamma}_{a}$, $\boldsymbol{\beta}_{a}$ are the scale and shift for query point $a$, respectively. Part (iii) of Figure 2 demonstrates that the kernel modulation enables query points, despite having identical receptive fields and translation vectors of input points, to output different kernel weights.
@@ 4.2 Neural Architecture for Discretization-Agnostic Reduced-Order PDE-Solving
The architecture of the CALM-PDE model follows an encode-process-decode paradigm (Sanchez-Gonzalez et al., 2020). Figure 2 shows an overview of the CALM-PDE architecture.
@@
[…]
@@
$d_{\theta}(Z^{t},P,\{\omega_{j}\}_{j=1}^{N^{\prime}})=l_{k}\Bigl{(}l_{k-1}\bigl{(}\dots l_{1}(Z^{t},P,A_{1}),A_{k-2},A_{k-1}\bigr{)},A_{k-1},\{\omega_{j}\}_{j=1}^{N^{\prime}}\biggr{)}=\{u^{t}(\omega_{j})\}_{j=1}^{N^{\prime}}$ (9)
@@ 5 Related Work
We give an overview of surrogate models, introduce continuous convolutional neural networks, and present reduced-order PDE-solving models. We refer to Appendix F for details on related work.
@@
Neural surrogate models are increasingly used for approximating the solutions of PDEs (Lu et al., 2019; Li et al., 2020a, b; Cao, 2021; Li et al., 2023a; Hagnberger et al., 2024). Neural operators (Kovachki et al., 2021) constitute an important category, and prevalent neural operator models include the Graph Neural Operator (GNO; Li et al. (2020b)) based on message passing (Gilmer et al., 2017) and the Fourier Neural Operator (FNO; Li et al. (2020a)), which leverages Fourier transforms, alongside its various variants. Transformers are also increasingly employed for surrogate modeling, typically using the attention mechanism on the spatial dimension of the PDE (Cao, 2021; Li et al., 2023a; Hao et al., 2023; Li et al., 2023b). Another distinct group of models, neural fields, is particularly well-suited for solving PDEs due to their ability to represent spatial functions. These models are gaining traction in PDE-solving applications (Yin et al., 2023; Chen et al., 2023; Serrano et al., 2023; Hagnberger et al., 2024; Knigge et al., 2024).
@@
6 Experiments and Evaluation

### v2

@@ 1 Introduction
Many scientific problems, such as climate modeling and fluid mechanics, rely on simulating physical systems, often involving solving spatio-temporal Partial Differential Equations (PDEs). In recent years, Machine Learning (ML) models have been successfully used to approximate the solution of PDEs (Lu et al., 2019; Li et al., 2020; Brandstetter et al., 2022; Chen & Wu, 2024), offering several advantages over classical numerical PDE solvers. For instance, ML models offer a data-driven approach that is applicable even if the underlying physics is (partially) unknown, can generate solutions more efficiently (Li et al., 2021b; Tompson et al., 2017), and are inherently differentiable by design, which is often not the case for numerical solvers (Takamoto et al., 2022).
@@
Practical applications usually require a densely discretized spatial domain $\Omega$, leading to more than 1M spatial points per timestep. For example, ML-based weather forecasting models typically operate on a spatial domain of $720\times 1440$ points or pixels (Pathak et al., 2022). Learning the solution of a time-dependent PDE directly in the physical domain $\Omega$ rather than in a compressed latent space can result in high memory and computational costs. Consequently, several architectures for reduced-order PDE-solving have been introduced (Wu et al., 2022; Alkin et al., 2024; Serrano et al., 2024; Wang & Wang, 2024). These models adopt the encode-process-decode paradigm (Sanchez-Gonzalez et al., 2020), wherein the physical domain is encoded into a compact latent space, its dynamics are evolved via a processor, and the output is decoded back to the original domain. Many real-world applications involve geometries with irregularly sampled spatial domains, necessitating discretization-agnostic models to effectively process such data. Unfortunately, existing approaches mainly utilize Convolutional Neural Networks (CNNs), which necessitate a regular spatial discretization (Wu et al., 2022), or memory-intensive attention-based mechanisms (Vaswani et al., 2017) to enable the encoding and decoding of arbitrarily discretized spatial domains (Alkin et al., 2024; Serrano et al., 2024; Wang & Wang, 2024).
@@
Motivated by these considerations, we introduce CALM-PDE (Continuous and Adaptive Convolutions for Latent Space Modeling of Time-dependent PDEs), an architecture featuring a novel encoder and decoder for reduced-order modeling of arbitrarily discretized PDEs. CALM-PDE compresses the spatial domain into a fixed latent representation with an encoder that builds on parametric continuous convolutional neural networks (Wang et al., 2018) and adaptively learns where to apply the convolution operator (learnable query points). This enables CALM-PDE to selectively sample more densely in important regions of the spatial domain, such as near complex solid boundaries, while allocating fewer points to smoother regions. Furthermore, we incorporate a locality inductive bias into the kernel function, consisting of an epsilon neighborhood and a distance-weighting, to enhance computational efficiency and to facilitate the learning of local patterns. An autoregressive model computes the temporal evolution completely in the latent space (latent time-stepping), which can be interpreted as Neural Ordinary Differential Equation (NODE; Chen et al. (2018)). The discretization-agnostic decoder, which builds on similar layers as the encoder, enables querying the output solution at arbitrary spatial locations. An overview of the CALM-PDE framework is shown in Figure˜1.
@@
  • We propose a model featuring a novel encoder and decoder that solves arbitrarily discretized PDEs in a fixed latent space.
@@ 2 Problem Definition
In the following section, we formally introduce the problem of solving PDEs. We refer to Appendix˜B for information about the dataset and training objective to train neural surrogates for PDE-solving.
@@
where $\boldsymbol{u}:[0,T]\times\Omega\rightarrow\mathbb{R}^{N_{c}}$ represents the solution function of the PDE that satisfies the Initial Condition (IC) $\boldsymbol{u}(0,\boldsymbol{\omega})$ for $t=0$ and the boundary conditions $\mathcal{B}[\boldsymbol{u}](t,\boldsymbol{\omega})$ if $\boldsymbol{\omega}$ is on the boundary $\partial\Omega$ of the domain $\Omega$. $N_{c}$ denotes the number of output channels or field variables of the PDE. Solving a PDE involves computing (an approximation of) the function $u$ that satisfies Equation˜1.
@@ 3 Background and Preliminaries
We briefly introduce deep parametric continuous convolutional neural networks (Wang et al., 2018). For comparison, discrete convolution and discrete CNNs are explained in Appendix˜C.
@@
Let $f,k:\mathbb{R}\rightarrow\mathbb{R}$ be two real-valued functions and $a\in\mathbb{R}$, then $(f*k)(a):=\int\displaylimits_{-\infty}^{\infty}f(\alpha)\cdot k(a-\alpha)d\alpha$ is the convolution of $f$ and $k$. Wang et al. (2018) propose to approximate the function $k$, which is the learnable kernel, with a Multi-Layer Perceptron (MLP) and the integral with Monte Carlo integration, which yields
$(f*k)(a)=\int\displaylimits_{-\infty}^{\infty}f(\alpha)\cdot k(a-\alpha)d\alpha\approx\frac{1}{N}\sum_{n=1}^{N}f(\alpha_{n})\cdot k(a-\alpha_{n})$ (2)
@@
where $\mathtt{RF}(a_{j})$ outputs the indices for the $M$ points that lie in the receptive field of a query point $a_{j}$. The receptive field can be constructed by only considering the K-nearest neighbors of the point $a_{j}$ or by considering the points in an epsilon neighborhood. Setting $M:=N$ means that the receptive field is not limited. Appendix˜D shows a visualization of continuous convolution. To generalize the layer from 1D to d-D, only the input dimension of the kernel function has to be adapted.
@@ 4.1 Learning Continuous Convolutions with Adaptive Query Points
We follow the formulation of a deep parametric continuous convolutional layer in Equation˜4 introduced by Wang et al. (2018) and propose the following improvements for encoding and decoding PDE solutions, which yield the CALM layer used for the encoder and decoder.
@@
We use a 2-layer MLP that takes $a-\alpha$ as input to parameterize the continuous kernel function $k_{i,o}(a-\alpha)$. The translation $a-\alpha$ is encoded with Random Fourier Features ($\mathtt{RFF}$; Li et al. (2021a)) to allow the kernel function to be less smooth (i.e., a small change in the input translation can lead to large output changes in the weight) which could be beneficial for high-frequency content where slightly different translation vectors should result in a large difference. Thus, the parametrized kernel function is given as $k_{i,o}(a-\alpha)=\mathtt{MLP}\bigl(\mathtt{RFF}(a-\alpha)\bigl)_{i,o}$ where $i$ is the $i^{th}$ input channel and $o$ the $o^{th}$ output channel.
@@
$k_{i,o}(a-\alpha)=\mathtt{MLP}\bigl(\mathtt{RFF}(a-\alpha)\bigr)_{i,o}\frac{\exp\left(-\frac{\left\lVert a-\alpha\right\rVert^{2}-\min(a)}{T(\max(a)-\min(a))}\right)}{\sum\limits_{\alpha_{j}\in\mathtt{RF}(a)}\exp\left(-\frac{\left\lVert a-\alpha_{j}\right\rVert^{2}-\min(a)}{T(\max(a)-\min(a))}\right)}$ (5)
for the kernel function with $\min(a)=\min\limits_{\alpha_{j}\in\mathtt{RF}(a)}\left\lVert a-\alpha_{j}\right\rVert^{2}$. This emphasizes input points closer to the query point within the epsilon neighborhood. The $\mathtt{softmax}$ also serves as a normalization factor for the Monte Carlo integration. The combination of the limited receptive field and distance weighting is a locality inductive bias, which is helpful because local patterns are often more important in PDE-solving (Chen & Wu, 2024). Part (i) of Figure˜2 shows the receptive field and the distance weighting within the receptive field.
@@
In continuous convolution, the input and output points do not have to be the same. This means that downsampling and upsampling are inherently supported, in contrast to CNNs, where strided convolution and variants are needed for downsampling. This allows us to reduce the number of points for encoding and to increase the number of points for decoding. We propose to make the query points for both encoding and decoding learnable. This allows the model to sample more densely in regions with important characteristics (e.g., a region that contains turbulence or an obstacle). The query positions are initially uniformly sampled from the domain $\Omega$. We assume that $\Omega$ is normalized in a range $[0,1)$ which yields $A=\{a_{j}\}_{j=1}^{K}=\{(x_{j},y_{j})\}_{j=1}^{K}$ with $x_{j},y_{j}\sim\mathcal{U}(0,1)$ for the initialization of the query points in 2D. Alternatively, the query points can be initialized by sampling $K$ query points from the underlying mesh. We call this initialization method of the query points as “mesh prior”. During the training, the query points are moved with a variant of stochastic gradient descent such as Adam (Kingma & Ba, 2017) as $(x_{j},y_{j})^{\top}\leftarrow(x_{j},y_{j})^{\top}-\eta\frac{\partial\mathcal{L}_{\boldsymbol{\theta}}}{\partial(x_{j},y_{j})^{\top}}\quad\forall j$ where $\eta$ denotes the learning rate and $\mathcal{L}_{\boldsymbol{\theta}}$ the cost function. The kernel function provides feedback in which direction to move, and the $\mathtt{softmax}$ function in Equation˜5 avoids hard cuts caused by the learnable query points and epsilon neighborhood by weighting the input points on the boundary of the epsilon neighborhood lower. The learnable query points are illustrated in part (ii) of Figure˜2.
@@
We allow each query point to have a customized filter kernel (i.e., different kernel weights for the same or similar translation vectors). This way, there can be multiple query points at the same location, each paying attention to different features. This could be helpful for simulations that contain a geometry or obstacles (e.g., many query points close to the obstacle, but each query learns different characteristics). The kernel modulation is done by scaling and shifting the intermediate representations (Perez et al., 2018) of the MLP in Equation˜5. Each query point has a location as well as scaling and shifting parameters. We denote the modulated kernel function as
@@
where we exclude the distance weighting for the sake of simplicity. $\boldsymbol{W}$ denotes a weight matrix with a suitable shape, $\sigma$ is a non-linearity, and $\boldsymbol{\gamma}_{a}$, $\boldsymbol{\beta}_{a}$ are the scale and shift for query point $a$, respectively. Part (iii) of Figure˜2 demonstrates that the kernel modulation enables query points, despite having identical receptive fields and translation vectors of input points, to output different kernel weights.
@@ 4.2 Neural Architecture for Discretization-Agnostic Reduced-Order PDE-Solving
The architecture of the CALM-PDE model follows an encode-process-decode paradigm (Sanchez-Gonzalez et al., 2020). Figure˜2 shows an overview of the CALM-PDE architecture.
@@
The model evolves the dynamics within the latent space through an autoregressive prediction. In particular, the processor $\psi_{\theta}$ predicts the difference from the latent tokens $Z^{t}$ at time $t$ to the future latent tokens $Z^{t+\Delta t}=\Delta t\cdot\psi_{\theta}(Z^{t},P)+Z^{t}$. To obtain the solution for a timestep $t+n\cdot\Delta t$, the processor is applied $n$ times iteratively to completely evolve the dynamics in the latent space (latent time-stepping). We opt for a Transformer model (Vaswani et al., 2017) since attention allows the processor to capture global information from local tokens, and latent tokens can dynamically interact with each other. For instance, the processor can learn the dynamics of a moving vortex in a CFD simulation by adapting the latent tokens on the path of the moving vortex and pushing it from one position to another. We include the positional information of the latent tokens into self-attention using the Euclidean distance, similar to Chen & Wu (2024). We refer to Section˜E.2 for details. Using the processor to predict the residual and scaling it with the temporal resolution $\Delta t$ can be understood as solving a NODE parametrized by $\psi_{\theta}$, which is given as $Z(t+\Delta t)=Z(t)+\int_{t}^{t+\Delta t}\psi_{\theta}(Z(\tau),P)d\tau$ with an explicit Euler solver. Thus, the processor implicitly defines a vector field that could be integrated with solvers more advanced than the explicit Euler method, such as a Runge-Kutta solver.
@@
$d_{\theta}(Z^{t},P,\{\omega_{j}\}_{j=1}^{N^{\prime}})$ $=l_{k}\Bigl(l_{k-1}\bigl(\dots l_{1}(Z^{t},P,A_{1}),A_{k-2},A_{k-1}\bigr),A_{k-1},\{\omega_{j}\}_{j=1}^{N^{\prime}}\biggr)$ (9)
$=\{u^{t}(\omega_{j})\}_{j=1}^{N^{\prime}}$
@@ 5 Related Work
We give an overview of surrogate models, introduce continuous convolutional neural networks, and present reduced-order PDE-solving models. We refer to Appendix˜F for details on related work.
@@
Neural surrogate models are increasingly used for approximating the solutions of PDEs (Lu et al., 2019; Li et al., 2021b, 2020; Cao, 2021; Li et al., 2023a; Hagnberger et al., 2024). Neural operators (Kovachki et al., 2023) constitute an important category, and prevalent neural operator models include the Graph Neural Operator (GNO; Li et al. (2020)) based on message passing (Gilmer et al., 2017) and the Fourier Neural Operator (FNO; Li et al. (2021b)), which leverages Fourier transforms, alongside its various variants. Transformers are also increasingly employed for surrogate modeling, typically using the attention mechanism on the spatial dimension of the PDE (Cao, 2021; Li et al., 2023a; Hao et al., 2023; Li et al., 2023b). Another distinct group of models, neural fields, is particularly well-suited for solving PDEs due to their ability to represent spatial functions. These models are gaining traction in PDE-solving applications (Yin et al., 2023; Chen et al., 2023; Serrano et al., 2023; Hagnberger et al., 2024; Knigge et al., 2024).
@@ 6 Experiments and Evaluation
6.1 Datasets
For regularly sampled spatial domains, we conduct experiments on the 1D Burgers’ equation dataset from PDEBench (Takamoto et al., 2022), the 2D Navier-Stokes equation datasets introduced by Li et al. (2021b), and the 3D compressible Navier-Stokes dataset of Takamoto et al. (2022). The 2D Euler equation with an airfoil geometry and the incompressible Navier-Stokes equation with cylinder geometries datasets from Pfaff et al. (2020) are used to evaluate the models on irregularly sampled domains. We refer to Appendix˜H for additional details on the datasets.
6.2 Baseline Models
We compare the CALM-PDE model against FNO (Li et al., 2021b) and Geo-FNO (Li et al., 2023c) for irregularly sampled meshes, F-FNO (Tran et al., 2023), OFormer (Li et al., 2023a), PIT (Chen & Wu, 2024), LNO (Wang & Wang, 2024), and AROMA (Serrano et al., 2024). We train the autoregressive models (FNO, F-FNO, Geo-FNO, and PIT) with a curriculum strategy that slowly increases the trajectory length, which improves the error compared to training with a full autoregressive rollout (see Appendix˜I). OFormer uses a similar strategy, LNO a one-step training, and AROMA a denoising diffusion training to learn the temporal dynamics of the PDE as proposed by the authors.
6.3 Results
We report the mean values and percentage deviations of the relative L2 error of multiple runs with different initializations. We opt for the relative L2 error as an evaluation metric because it weights channels with small and large magnitudes equally and does not ignore time-dependent decay effects like in the 1D Burgers’ equation. We refer to Section˜N.1 for the full results with standard deviations and Section˜N.2 for qualitative results.
RQ1.
We train and evaluate the models on regularly sampled spatial domains. Regularly sampled spatial domains refer to meshes with equidistant nodes or points. We select datasets that span the entire spatial dimension from 1D to 3D. Table˜1 shows the relative L2 errors of the baselines and CALM-PDE model. On 2 out of 4 benchmark problems, CALM-PDE outperforms all baselines. On one problem, CALM-PDE achieves the third-lowest errors and on the remaining benchmark problem, CALM-PDE is outperformed by F-FNO and LNO, and outperforms PIT and OFormer.